If a normal chord of y^2=4ax subtends an right angle at the vertex of the parabola, then it's slope is equal to?
A)  +-1
 B) +-root2
 C) +-2
D)  None

Grade:

1 Answers

kkbisht
5 years ago
the answer is ± \sqrt{}2
explanation:  Let the normal chord at A( at2,2at) meet the parabola at the point B(at21,2at1) and since \angle AOB=\pi/2
therefore tt1=-4 and also the condition for t1= -t -2/t
therefore we have  -4/t =-t -2/t ( using tt1=-4 ) => 2/t=t => t2=2 => t= +-\sqrt{}2
 
kkbisht